# A Comprehensive Survey on Knowledge Distillation

### v1

@@ Abstract
Index Terms:
Knowledge Distillation, Knowledge Transfer, Teacher-Student Architecture.
I Introduction
With the emergence of DNNs, the fields of Computer Vision (CV) and Natural Language Processing (NLP) have been revolutionized, and most tasks in these domains are now solved using DNNs. While a simple ResNet [1] model or BERT [2] can be easily trained on most of existing GPUs, the advent of large models like LLMs and foundational vision models has made training and inference a significant challenge in terms of runtime and memory usage, especially for deployment on edge devices like mobile phones due to their widespread use. Despite their high performance, these large models often have complex architectures and are heavily overparameterized [3, 4]. For example, the weight matrices in LLMs have been shown to be low-rank matrices [5].
@@ I Introduction
To address the issue of large models, various solutions have been proposed, such as efficient network architectures, compression methods, pruning, quantization, low-rank factorization, and knowledge distillation. Efficient network blocks, including MobileNet [6], ShuffleNet [7], and BiSeNet [8] have been introduced in recent years. Pruning methods, a type of compression technique, aim to remove unnecessary layers and parameters from models with minimal impact on performance. Low-rank factorization methods reduce parameters through matrix decomposition. Unlike other methods, KD does not modify the network’s layers or parameters. In KD, a lightweight network (student) is trained under the supervision of a more complex model (teacher) with deeper architecture and greater number of parameters. This concept, first proposed by [9] and later popularized by [10] as KD, introduced the idea of training the student network by mimicking the teacher’s output distribution as soft labels.
@@
With the significant growth in the number of published papers in the field of KD and its wide applications across various tasks and domains, several survey papers have been presented to review KD from different perspectives. The prominent work [11] provides a comprehensive survey on KD, reviewing it from the aspects of knowledge types, algorithms, and schemes, while including comparisons between different methods. Another notable work, [12], offers an overview of model compression based on the teacher-student architecture specifically for vision tasks. [13] presents a survey on KD and proposes a new metric for comparing distillation methods based on size and accuracy, while [14] categorizes existing approaches based on the type of knowledge source used for distillation. More recently, [15] delivers a comprehensive survey, exploring knowledge optimization objectives associated with various representations, and [16] updates earlier surveys by providing an extensive review of KD methods. Additionally, they briefly investigate distillation in VLMs and discuss the challenges of distillation under limited data scenarios.
While each of these surveys provides a detailed summary of papers from different perspectives, they also have some drawbacks. First, although these surveys analyze existing approaches in each category, they fail to address recent advancements, particularly in feature-based distillation methods. Despite the significant growth of feature-based distillation, which now dominates state-of-the-art methods, many recent and impactful works are overlooked. Furthermore, adaptive distillation and contrastive distillation algorithms are rarely discussed. Second, with the recent emergence of foundation models and LLMs, these surveys have largely overlooked their potential for distilling knowledge into other models. One significant shortcomings of [11], as the most prominent survey in the field, is its lack of explanation regarding the applications of distillation in newly introduced models such as foundation models and LLMs, which have gained significant attention in recent years. Third, none of the existing surveys have investigated KD for 3D inputs like point clouds. As 3D tasks garner increasing attention in top research venues, the lack of sufficient approaches and models in comparison to the image domain highlights the importance of KD as an effective method for training the models in this area. Table I shows a summarized comparison between existing surveys and this work.
TABLE I: Comparison between existing KD surveys and this survey.
@@
Gou et al. [11] Logit, Feature, Similarity Adversarial, Attention, Cross-modal, Data-free, Graph, Lifelong, Multi-teacher, NAS, Quantized Image, Text, Speech, Video Visual Recognition
Wang et al. [12] Logit, Feature, Similarity Adversarial, Cross-modal, Data-free, Few-shot, Graph, Incremental, Multi-teacher, Reinforcement Image, Video Visual Recognition, Self-supervised Learning
Hu et al. [15] Logit, Feature, Similarity, Mutual Information Cross-modal, Federated, Graph, Multi-teacher Image, Text, Speech, Video Visual Recognition, Ranking, Generation, Regression
Moslemi et al. [16] Logit, Feature, Similarity Adversarial, Attention, Cross-modal, Data-free, Graph, Lifelong, Multi-teacher, NAS, Quantized Image, Text Vision-Language Models, Medical Image
@@
II Sources
The primary component of a distillation method is the source of knowledge being distilled. Various sources of knowledge can be utilized for distillation. In the initial concept of distillation [10], logits were employed as the teacher’s final output. Alongside logit distillation [17, 18, 19, 20], the activations [21], neurons, and features [22, 23] of intermediate layers play a crucial role in guiding the student network. Moreover, relationships and similarities among channels [24], instances [25], and classes [26] provide higher-order information that facilitates the transfer of knowledge from the teacher to the student. In this section, KD methods based on their sources of distillation are examined. Specifically, existing methods are classified into three categories: Logit-based, Feature-based, and Similarity-based as shown in Figure 3. Table II summarizes logit-based and similarity-based distillation methods and Table III presents detailed information on existing feature distillation methods. Subsequently, each type of distillation is explained in detail.
@@ II Sources
II-A Logit-based Distillation
The most straightforward source for KD is logits. Logits are outputs of the last layer of the model, and the idea is to force the student model to mimic the final predictions of the teacher, which are supposed to contain informative dark knowledge [10] of the teacher. The prediction from the last layer can vary for each task. For example, logits in classification contain predictions for each class, in detection they contain predicted coordinates, and in pose estimation, they can represent heatmaps. In general, let $Z_{T}$ and $Z_{S}$ be logits of the last layers of the teacher and student, respectively. Then the logit-based distillation loss ($L_{ld}$) is defined as
@@ II-A Logit-based Distillation
The first papers on KD were proposed for image classification, where soft-labels were introduced by applying softmax on logits [10, 17]. These output distributions, or soft-labels, provide information on the probability of the input belonging to each class. The distillation loss function for soft-labels is expressed as
@@
where $exp$ is softmax function, $\tau$ is the temperature factor that controls the softening of logits, and $KL(\cdot)$ is the Kullback–Leibler Divergence (KL) function. By optimizing the above objective, the student can mimic the output distribution of the teacher and make better predictions.
@@
Although the initial concept of KD was simple and effective, many different logit-based methods have been proposed in recent years [18, 27, 19, 28]. These methods mainly aim to normalize logits before distillation [29, 30, 31, 32, 20], soften or smoothen the logits [28, 33, 34], or decouple the logit-distillation loss [35, 36, 37, 38, 39, 40]. [29] finds that the magnitude of confidence is not necessary for KD and proposes spherical KD to reduce the gap between teacher and student. [30] shows that it is not feasible to soften all the samples with a constant temperature and proposes NormKD to customize the temperature for each sample according to the characteristics of the sample’s logit distribution. [31] and [32] investigate the role of projectors and temperature in the KD process respectively. [20] suggests logit standardization by setting the temperature as the weighted standard deviation of the logit and performing a plug-and-play Z-score pre-process. On the other hand, among methods that try to soften the logits, [28] perturbs the logits for distillation with some noises, and [34] softens the logits before KD and utilizes a learning simplifier with an attention module. However, most existing methods aim to decouple the distillation loss. For instance, [35, 36, 37] separate the distillation of target-class and non-target class distributions. [39] reshapes the logits for multi-scale logit distillation, [40] uses different weights for distilling edges and bodies of objects, and [38] decouples the KL to a weighted MSE and a cross-entropy loss that incorporates soft-labels.
@@
In summary, logit distillation is a simple and straightforward method that can be likened to label smoothing [41] and regularization [42, 43], aiming to enhance the student network’s performance by preventing overfitting [11]. However, due to the reliance of logits on predefined classes, its application is limited to supervised learning. Moreover, mimicking the last-layer predictions does not provide insights into the intermediate layers of the teacher network. It can be challenging for the student to learn effectively solely from the final outputs of the teacher, especially when teacher and student networks have different architectures. Therefore, in addition to logit distillation, there is a necessity to distill intermediate outputs and similarities from the teacher network to facilitate effective knowledge transfer.
TABLE II: Summary of logit-based and similarity-based distillation methods.
@@
Logit Logit Normalization Standardizes the logits before distillation KD [10], SphericalKD [29], NormKD [30], Miles et al. [31], TTM [32], Sun et al. [20]
Logit Softening Smooths the logits by adding noise Sau et al. [28], SFKD [34], Li et al. [33]
Decoupling Splits the logit distillation loss into separate terms DKL [38], BPKD [40], SDD [39], Ding et al. [37], NTCE [36]
Revisiting Logit Distillation Reduces the gap between teacher and student TAKD [18], VanillaKD [19], EffDstl [27]
Similarity Instance-level Similarity Pairwise similarity between different data samples MTKD [44], RKD [45], SPKD [25], IRGKD [46], Chen et al. [47], CSKD [48], Wang et al. [49], Cheng et al. [50], GIRKD [51]
Feature/Channel-level Similarity Pairwise similarity between features/channels Yim et al. [52], ICKD [24], CCD [53]
Class-level Similarity Pairwise similarity between classes CSKD [48], DSD [54], DistKD [55], IDD [56], AICSD [57]
Others Similarity between regions/pixels/neighbors/views SKD [58], CIRKD [59], PRRD [60], BCKD [61], NRKD [62], CRLD [63]
II-B Feature-based Distillation
@@
The concept of feature distillation was first explored by [22], where they proposed providing the student network with hints from the teacher. More specifically, they suggested minimizing the differences in features between the teacher and student to align them. Generally, let $F_{s}$ and $F_{t}$ be intermediate features of the teacher and student, respectively. A general feature distillation loss ($L_{fd}$) between teacher and student can then be defined as follows
@@
Inspired by the idea of FitNet, several feature-based distillation methods have been proposed. Initially, AT[23] introduced attention transfer by minimizing the $L_{2}$ norm of the difference between the feature maps of the teacher and student. [64] focused on matching the distributions of neuron selectivity patterns between the teacher and student by minimizing the Maximum Mean Discrepancy (MMD). [21] transferred the activation boundaries formed by hidden neurons, while [65] transferred the factors of the features. [66] presented a comprehensive overhaul of feature distillation, proposing to distill the pre-activation feature maps, and [67] introduced channel-wise distillation by applying softmax to channels of features and matching the distributions between teacher and student.
@@
Subsequently, some methods explored cross-layer feature distillation. [68] introduced Knowledge Review, where the feature maps of each layer of the teacher should match the student’s feature maps in the corresponding layer and all features from the previous layers. [69] proposed a more general framework where each feature from the student should match all of the teacher’s features, learning weights to determine the connection between the feature maps. This framework is a generalized version of FitNet, where the weights for corresponding features are set to one.
@@
Recently, the focus of many papers has shifted towards the transformation function of the features, highlighting its role in feature matching beyond just matching feature sizes. [21] trains an autoencoder to transform the student’s features for better alignment with the teacher’s features. [70] randomly masks some pixels of the student’s feature map and uses two convolution layers to transform the masked features, aligning them with the corresponding features in the teacher’s model to encourage the student to produce feature maps similar to those of the teacher. [71] demonstrated that masking is not necessary and a simple Multi-Layer Perceptron (MLP) layer for transforming the student’s features is sufficient.
@@
Most recently, [72] combined the idea of diffusion [73] with KD, training a diffusion model on the teacher’s feature maps and using it to denoise the student’s feature maps. It considers the student’s features as a noisy version of the teacher’s features and employs diffusion as a transformation for the student’s features. In a concurrent work, [74] used Convolutional Block Attention Mechanism (CBAM) [75] as a transformation, applying channel and spatial attention to the teacher’s and student’s features to highlight the most discriminative parts. It defines an MSE loss between the refined feature maps of the teacher and student.
@@
Concurrent with the research path outlined, several other methods have been proposed. [76] rethinks raw feature alignment and decomposes the loss function of FitNets [22] into a magnitude difference term and an angular difference term. [77] matches the features of the teacher with augmented versions of the student’s features, while [78] distills features in the frequency domain. [79] proposes N-to-one matching between the features of the student and teacher, respectively, and [80] suggests that distilling corresponding features results in the disregard of some layers of the teacher. Instead, they propose to fuse all the features of the teacher for distillation.
@@
TABLE III: Summary of feature distillation methods.
@@
FitNet [22] (2014) $1\times 1$ Conv Feature Representation $\mathcal{L}_{2}(\cdot)$
AT [23] (2016) Channel Aggregation Channel Aggregation Attention Map $\mathcal{L}_{2}(\cdot)$
NST [64] (2017) $1\times 1$ Conv Neuron Selectivity Pattern $\mathcal{L}_{MMD}(\cdot)$
Jacobian [81] (2018) Gradient Gradient Jacobian of Feature $\mathcal{L}_{2}(\cdot)$
FT [65] (2018) Auto-encoder Auto-encoder Paraphraser $\mathcal{L}_{1}(\cdot)$
AB [21] (2019) Binarization $1\times 1$ Conv Activation Boundary Marginal $\mathcal{L}_{2}$
Heo et al. [66] (2019) Margin ReLU $1\times 1$ Conv Pre-ReLU Feature Partial $\mathcal{L}_{2}$
He el al. [82] (2019) Auto-encoder $3\times 3$ Conv Adapted Feature $\mathcal{L}_{1}(\cdot)$/$\mathcal{L}_{2}(\cdot)$
FN [83] (2020) $L_{2}$ Normalization $L_{2}$ Normalization Feature Representation $\mathcal{L}_{CE}(\cdot)$
CWD [67] (2021) $\text{Softmax}(\cdot)$ $\text{Softmax}(\cdot)$ Activation Map $KL(\cdot)$
MGD [70] (2022) Masking + Conv Feature Representation $\mathcal{L}_{2}(\cdot)$
MLP [71] (2023) MLP Feature Representation $\mathcal{L}_{2}(\cdot)$
DiffKD [72] (2023) Diffusion Process Denoised Feature $KL(\cdot)$/$\mathcal{L}_{2}(\cdot)$
FAKD [77] (2024) Feature Augmentation Activation Boundary $KL(\cdot)$
LAD [76] (2024) $1\times 1$ Conv Angular Information of Feature $\mathcal{L}_{2}(\cdot)$
AttnFD [74] (2024) Channel/Spatial Attention Channel/Spatial Attention Refined Feature $\mathcal{L}_{2}(\cdot)$
II-C Similarity-based Distillation

### v2

@@ Abstract
1 Introduction
With the emergence of DNNs, the fields of Computer Vision (CV) and Natural Language Processing (NLP) have been revolutionized, and most tasks in these domains are now solved using DNNs. While a simple ResNet (He et al., 2016) model or BERT (Kenton & Toutanova, 2019) can be easily trained on most of existing GPUs, the advent of large models like LLMs and foundational vision models has made training and inference a significant challenge in terms of runtime and memory usage, especially for deployment on edge devices like mobile phones due to their widespread use. Despite their high performance, these large models often have complex architectures and are heavily over-parameterized (Kaplan et al., 2020; Fischer et al., 2024). For example, the weight matrices in LLMs have been shown to be low-rank matrices (Hu et al., 2021).
@@ 1 Introduction
To address the issue of large models, various solutions have been proposed, such as efficient network architectures, compression methods, pruning, quantization, low-rank factorization, and knowledge distillation. Efficient network blocks, including MobileNet (Howard, 2017), ShuffleNet (Zhang et al., 2018), and BiSeNet (Yu et al., 2018) have been introduced in recent years. Pruning methods, a type of compression technique, aim to remove unnecessary layers and parameters from models with minimal impact on performance. Low-rank factorization methods reduce parameters through matrix decomposition. Unlike other methods, KD does not modify the network’s layers or parameters. In KD, a lightweight network (student) is trained under the supervision of a more complex model (teacher) with deeper architecture and greater number of parameters. This concept, first proposed by Buciluǎ et al. (2006) and later popularized by Hinton (2015) as KD, introduced the idea of training the student network by mimicking the teacher’s output distribution as soft labels.
@@
With the significant growth in the number of published papers in the field of KD and its wide applications across various tasks and domains, several survey papers have been presented to review KD from different perspectives. The prominent work Gou et al. (2021) provides a comprehensive survey on KD, reviewing it from the aspects of knowledge types, algorithms, and schemes, while including comparisons between different methods. Another notable work, Wang & Yoon (2021), offers an overview of model compression based on the teacher-student architecture specifically for vision tasks. Alkhulaifi et al. (2021) presents a survey on KD and proposes a new metric for comparing distillation methods based on size and accuracy, while Yang et al. (2023c) categorizes existing approaches based on the type of knowledge source used for distillation. More recently, Hu et al. (2023a) delivers a comprehensive survey, exploring knowledge optimization objectives associated with various representations, and Moslemi et al. (2024) updates earlier surveys by providing an extensive review of KD methods. Additionally, they briefly investigate distillation in VLMs and discuss the challenges of distillation under limited data scenarios.
While each of these surveys provides a detailed summary of papers from different perspectives, they also have some drawbacks. First, although these surveys analyze existing approaches in each category, they fail to address recent advancements, particularly in feature-based distillation methods. Despite the significant growth of feature-based distillation, which now dominates state-of-the-art methods, many recent and impactful works are overlooked. Furthermore, adaptive distillation and contrastive distillation algorithms are rarely discussed. Second, with the recent emergence of foundation models and LLMs, these surveys have largely overlooked their potential for distilling knowledge into other models. One significant shortcomings of Gou et al. (2021), as the most prominent survey in the field, is its lack of explanation regarding the applications of distillation in newly introduced models such as foundation models and LLMs, which have gained significant attention in recent years. Third, none of the existing surveys have investigated KD for 3D inputs like point clouds. As 3D tasks garner increasing attention in top research venues, the lack of sufficient approaches and models in comparison to the image domain highlights the importance of KD as an effective method for training the models in this area. Table 1 shows a summarized comparison between existing surveys and this work.
Table 1: Comparison between existing KD surveys and the current survey. New topics not covered in previous works are highlighted in bold to indicate the progress and scope of this survey.
@@
Gou et al. (2021) Logit, Feature, Similarity Adversarial, Attention, Cross-modal, Data-free, Graph, Lifelong, Multi-teacher, NAS, Quantized Image, Text, Speech, Video Visual Recognition
Wang & Yoon (2021) Logit, Feature, Similarity Adversarial, Cross-modal, Data-free, Few-shot, Graph, Incremental, Multi-teacher, Reinforcement Image, Video Visual Recognition, Self-supervised Learning
Hu et al. (2023a) Logit, Feature, Similarity, Mutual Information Cross-modal, Federated, Graph, Multi-teacher Image, Text, Speech, Video Visual Recognition, Ranking, Generation, Regression
Moslemi et al. (2024) Logit, Feature, Similarity Adversarial, Attention, Cross-modal, Data-free, Graph, Lifelong, Multi-teacher, NAS, Quantized Image, Text Vision-Language Models, Medical Image
@@
2 Sources
The primary component of a distillation method is the source of knowledge being distilled. Various sources of knowledge can be utilized for distillation. In the initial concept of distillation (Hinton, 2015), logits were employed as the teacher’s final output. Alongside logit distillation (Ba & Caruana, 2014; Mirzadeh et al., 2020; Hao et al., 2023; Sun et al., 2024a), the activations (Heo et al., 2019b), neurons, and features (Romero et al., 2014; Zagoruyko & Komodakis, 2016) of intermediate layers play a crucial role in guiding the student network. Moreover, relationships and similarities among channels (Liu et al., 2021c), instances (Tung & Mori, 2019), and classes (Wang et al., 2020d) provide higher-order information that facilitates the transfer of knowledge from the teacher to the student. In this section, KD methods based on their sources of distillation are examined. Specifically, existing methods are classified into three categories: Logit-based, Feature-based, and Similarity-based as shown in Figure 3. Table 2 summarizes logit-based and similarity-based distillation methods and Table 3 presents detailed information on existing feature distillation methods. Subsequently, each type of distillation is explained in detail.
@@ 2 Sources
2.1 Logit-based Distillation
The most straightforward source for KD is logits. Logits are outputs of the last layer of the model, and the idea is to force the student model to mimic the final predictions of the teacher, which are supposed to contain informative dark knowledge (Hinton, 2015) of the teacher. The prediction from the last layer can vary for each task. For example, logits in classification contain predictions for each class, in detection they contain predicted coordinates, and in pose estimation, they can represent heatmaps. In general, let $Z_{T}$ and $Z_{S}$ be logits of the last layers of the teacher and student, respectively. Then the logit-based distillation loss ($L_{ld}$) is defined as
@@ 2.1 Logit-based Distillation
The first papers on KD were proposed for image classification, where soft-labels were introduced by applying softmax on logits (Hinton, 2015; Ba & Caruana, 2014). These output distributions, or soft-labels, provide information on the probability of the input belonging to each class. The distillation loss function for soft-labels is expressed as
@@
where expexp is softmax function, $\tau$ is the temperature factor that controls the softening of logits, and $KL(\cdot)$ is the Kullback–Leibler Divergence (KL) function. By optimizing the above objective, the student can mimic the output distribution of the teacher and make better predictions.
@@
Although the initial concept of KD was simple and effective, many different logit-based methods have been proposed in recent years (Sau & Balasubramanian, 2016; Mirzadeh et al., 2020; Song et al., 2023b; Hao et al., 2023; Xu et al., 2023d; Lu et al., 2025). These methods mainly aim to normalize logits before distillation (Guo et al., 2020a; Chi et al., 2023; Miles & Mikolajczyk, 2024; Zheng & Yang, 2024; Sun et al., 2024a), soften or smoothen the logits (Sau & Balasubramanian, 2016; Li et al., 2023a; Yuan et al., 2024b), or decouple the logit-distillation loss (Zhao et al., 2022a; Li et al., 2024a; Ding et al., 2024; Cui et al., 2024; Wei et al., 2024b; Liu et al., 2024d). Guo et al. (2020a) finds that the magnitude of confidence is not necessary for KD and proposes spherical KD to reduce the gap between teacher and student. Chi et al. (2023) shows that it is not feasible to soften all the samples with a constant temperature and proposes NormKD to customize the temperature for each sample according to the characteristics of the sample’s logit distribution. Miles & Mikolajczyk (2024) and Zheng & Yang (2024) investigate the role of projectors and temperature in the KD process respectively. Sun et al. (2024a) suggests logit standardization by setting the temperature as the weighted standard deviation of the logit and performing a plug-and-play Z-score pre-process. On the other hand, among methods that try to soften the logits, Sau & Balasubramanian (2016) perturbs the logits for distillation with some noises, and Yuan et al. (2024b) softens the logits before KD and utilizes a learning simplifier with an attention module. However, most existing methods aim to decouple the distillation loss. For instance, Zhao et al. (2022a); Li et al. (2024a); Ding et al. (2024) separate the distillation of target-class and non-target class distributions. Wei et al. (2024b) reshapes the logits for multi-scale logit distillation, Liu et al. (2024d) uses different weights for distilling edges and bodies of objects, and Cui et al. (2024) decouples the KL to a weighted MSE and a cross-entropy loss that incorporates soft-labels.
@@
In summary, logit distillation is a simple and straightforward method that can be likened to label smoothing (Kim & Kim, 2017) and regularization (Müller et al., 2019; Ding et al., 2019), aiming to enhance the student network’s performance by preventing overfitting (Gou et al., 2021). However, due to the reliance of logits on predefined classes, its application is limited to supervised learning. Moreover, mimicking the last-layer predictions does not provide insights into the intermediate layers of the teacher network. It can be challenging for the student to learn effectively solely from the final outputs of the teacher, especially when teacher and student networks have different architectures. Therefore, in addition to logit distillation, there is a necessity to distill intermediate outputs and similarities from the teacher network to facilitate effective knowledge transfer.
Table 2: Summary of logit-based and similarity-based distillation methods.
@@
Logit Logit Normalization Standardizes the logits before distillation KD (Hinton, 2015), SphericalKD (Guo et al., 2020a), NormKD (Chi et al., 2023), Miles et al. (Miles & Mikolajczyk, 2024), TTM (Zheng & Yang, 2024), Sun et al. (Sun et al., 2024a)
Logit Softening Smooths the logits by adding noise Sau et al. (Sau & Balasubramanian, 2016), Li et al. (Li et al., 2023a), SFKD (Yuan et al., 2024b)
Decoupling Splits the logit distillation loss into separate terms DKL (Cui et al., 2024), BPKD (Liu et al., 2024d), SDD (Wei et al., 2024b), Ding et al. (Ding et al., 2024), NTCE (Li et al., 2024a)
Revisiting Logit Distillation Reduces the gap between teacher and student TAKD (Mirzadeh et al., 2020), VanillaKD (Hao et al., 2023), EffDstl (Song et al., 2023b), Xu et al. (Xu et al., 2023d), BDCKD (Lu et al., 2025)
Similarity Instance-level Similarity Pairwise similarity between different data samples MTKD (You et al., 2017), RKD (Park et al., 2019), SPKD (Tung & Mori, 2019), IRGKD (Liu et al., 2019f), Chen et al. (Chen et al., 2020b), CSKD (Chen et al., 2020f), Wang et al. (Wang et al., 2024g), Cheng et al. (Cheng et al., 2024a), GIRKD (Hu et al., 2024b)
Feature/Channel-level Similarity Pairwise similarity between features/channels Yim et al. (Yim et al., 2017), ICKD (Liu et al., 2021c), CCD (Wang et al., 2023d)
Class-level Similarity Pairwise similarity between classes CSKD (Chen et al., 2020f), DSD (Feng et al., 2021), DistKD (Huang et al., 2022a), IDD (Zhang et al., 2022d), AICSD (Mansourian et al., 2025), Dist+ (Huang et al., 2025a), IKD (Wang et al., 2025a)
Others Similarity between regions/ pixels/ neighbors/ views SKD (Liu et al., 2019d), CIRKD (Yang et al., 2022c), PRRD (Wang et al., 2023b), BCKD (Wang et al., 2023k), NRKD (Xin et al., 2024), CRLD (Zhang et al., 2024g)
2.2 Feature-based Distillation
@@
The concept of feature distillation was first explored by Romero et al. (2014), where they proposed providing the student network with hints from the teacher. More specifically, they suggested minimizing the differences in features between the teacher and student to align them. Generally, let $F_{s}$ and $F_{t}$ be intermediate features of the teacher and student, respectively. A general feature distillation loss ($L_{fd}$) between teacher and student can then be defined as follows
@@
Inspired by the idea of FitNet, several feature-based distillation methods have been proposed. Initially, AT(Zagoruyko & Komodakis, 2016) introduced attention transfer by minimizing the $L_{2}$ norm of the difference between the feature maps of the teacher and student. Huang & Wang (2017) focused on matching the distributions of neuron selectivity patterns between the teacher and student by minimizing the Maximum Mean Discrepancy (MMD). Heo et al. (2019b) transferred the activation boundaries formed by hidden neurons, while Kim et al. (2018) transferred the factors of the features. Heo et al. (2019a) presented a comprehensive overhaul of feature distillation, proposing to distill the pre-activation feature maps, and Shu et al. (2021) introduced channel-wise distillation by applying softmax to channels of features and matching the distributions between teacher and student.
@@
Subsequently, some methods explored cross-layer feature distillation. SimKD (Chen et al., 2022a) employed a reused teacher classifier. Chen et al. (2021d) introduced Knowledge Review, where the feature maps of each layer of the teacher should match the student’s feature maps in the corresponding layer and all features from the previous layers. Chen et al. (2021a) proposed a more general framework where each feature from the student should match all of the teacher’s features, learning weights to determine the connection between the feature maps. This framework is a generalized version of FitNet, where the weights for corresponding features are set to one.
@@
Recently, the focus of many papers has shifted towards the transformation function of the features, highlighting its role in feature matching beyond just matching feature sizes. Heo et al. (2019b) trains an autoencoder to transform the student’s features for better alignment with the teacher’s features. Yang et al. (2022f) randomly masks some pixels of the student’s feature map and uses two convolution layers to transform the masked features, aligning them with the corresponding features in the teacher’s model to encourage the student to produce feature maps similar to those of the teacher. Liu et al. (2023c) demonstrated that masking is not necessary and a simple Multi-Layer Perceptron (MLP) layer for transforming the student’s features is sufficient.
@@
Most recently, Huang et al. (2023b) combined the idea of diffusion (Ho et al., 2020) with KD, training a diffusion model on the teacher’s feature maps and using it to denoise the student’s feature maps. It considers the student’s features as a noisy version of the teacher’s features and employs diffusion as a transformation for the student’s features. In a concurrent work, Mansourian et al. (2024) used Convolutional Block Attention Mechanism (CBAM) (Woo et al., 2018) as a transformation, applying channel and spatial attention to the teacher’s and student’s features to highlight the most discriminative parts. It defines an MSE loss between the refined feature maps of the teacher and student.
@@
Concurrent with the research path outlined, several other methods have been proposed. Liu et al. (2024g) rethinks raw feature alignment and decomposes the loss function of FitNets (Romero et al., 2014) into a magnitude difference term and an angular difference term. Yuan et al. (2024a) matches the features of the teacher with augmented versions of the student’s features, while Zhang et al. (2024i) distills features in the frequency domain. Liu et al. (2023a) proposes N-to-one matching between the features of the student and teacher, respectively, and Passban et al. (2021) suggests that distilling corresponding features results in the disregard of some layers of the teacher. Instead, they propose to fuse all the features of the teacher for distillation.
@@
Table 3: Summary of feature distillation methods.
@@
FitNet (Romero et al., 2014) $1\times 1$ Conv Feature Representation $\mathcal{L}_{2}(\cdot)$
AT (Zagoruyko & Komodakis, 2016) Channel Aggregation Channel Aggregation Attention Map $\mathcal{L}_{2}(\cdot)$
NST (Huang & Wang, 2017) $1\times 1$ Conv Neuron Selectivity Pattern $\mathcal{L}_{MMD}(\cdot)$
Jacobian (Srinivas & Fleuret, 2018) Gradient Gradient Jacobian of Feature $\mathcal{L}_{2}(\cdot)$
FT (Kim et al., 2018) Auto-encoder Auto-encoder Paraphraser $\mathcal{L}_{1}(\cdot)$
AB (Heo et al., 2019b) Binarization $1\times 1$ Conv Activation Boundary Marginal $\mathcal{L}_{2}$
Heo et al. (Heo et al., 2019a) Margin ReLU $1\times 1$ Conv Pre-ReLU Feature Partial $\mathcal{L}_{2}$
He el al. (He et al., 2019) Auto-encoder $3\times 3$ Conv Adapted Feature $\mathcal{L}_{1}(\cdot)$$\mathcal{L}_{2}(\cdot)$
FN (Xu et al., 2020b) $L_{2}$ Normalization $L_{2}$ Normalization Feature Representation $\mathcal{L}_{CE}(\cdot)$
CWD (Shu et al., 2021) $\text{Softmax}(\cdot)$ $\text{Softmax}(\cdot)$ Activation Map $KL(\cdot)$
MGD (Yang et al., 2022f) Masking + Conv Feature Representation $\mathcal{L}_{2}(\cdot)$
MLP (Liu et al., 2023c) MLP Feature Representation $\mathcal{L}_{2}(\cdot)$
DiffKD (Huang et al., 2023b) Diffusion Process Denoised Feature $KL(\cdot)$$\mathcal{L}_{2}(\cdot)$
FAKD (Yuan et al., 2024a) Feature Augmentation Activation Boundary $KL(\cdot)$
LAD (Liu et al., 2024g) $1\times 1$ Conv Angular Information of Feature $\mathcal{L}_{2}(\cdot)$
AttnFD (Mansourian et al., 2024) Channel/Spatial Attention Channel/Spatial Attention Refined Feature $\mathcal{L}_{2}(\cdot)$
2.3 Similarity-based Distillation